OpenStudy (thomasfire55):

(x - 5)^2 = 81 Which of the following expressions represents the solutions to the given equation?

OpenStudy (thomasfire55):

a)-5 ± √(81) b)5 ± √(81) c)-25 ± √(81)

OpenStudy (mariamikayla):

(x-5)^2 =81 root from (x-5)^2 = root from 81 |x-5|= 9 x-5= 9 or x-5= -9 x=14 x= -4 or you can solve this in this method: (x-5)^2=81 x^2-10x+25-81=0 x^2-10x- 56=0 Delta= 100 -4*(-56) = 100+ 224= 324 x1;x2= (10 +/- 18)/2 x1= (10+18)/2= 28/2=14 x2=(10-18)/2= -8/2= -4

OpenStudy (mariamikayla):

You've written the variants right?

OpenStudy (thomasfire55):

?